# arXiv:2603.00704v1[stat.ME] 28 Feb 2026

### 28 Feb 2026 v1

@@ Abstract.
Version: . Details on all computations are available on request and will be made available on the web in due course.
@@ 1. Introduction
has been proposed as a plug-in estimator for $G$, Jiang and Zhang (2009), Koenker and Mizera (2014) and Soloff et al. (2024). This $G$-modeling strategy – in the terminology of Efron (2019) – performs well in simulations, e.g. Koenker and Mizera (2014), Gu and Koenker (2016), Koenker and Gu (2026), relative to alternatives that attempt to estimate $f_{G}$ directly or makes a priori assumptions about the form of $G$. However, it is obviously subject to the criticism that the Gaussian assumption on the likelihood is quite strong, and priors are never terribly precise.
@@ 3. Restricted Bayes Solutions
This is the problem posed by Mallows (1978) motivated by robustness considerations for time-series problems with additive outliers. Mallows conjectured that the least favorable $G$ would be discrete, supported on the integers with mass declining exponentially. Bickel (1983) reports a modified conjecture of Donoho that relaxes the spacing of the Mallows mass points, but is otherwise similar. Neither conjecture seems to be strictly correct, but numerical computations confirm the nearly exponential decay of the mass. Bickel and Collins (1983) provide a detailed discussion of the discrete nature of the Mallows solutions based on the analyticity of the objective function. Our Appendix A on computation of the Mallows least favorable distribution provides further details. See also Johnstone (1994).
@@
The Mallows rule has the obvious advantage that it yields a Bayes decision rule while the corresponding Huber procedure does not. This is particularly evident in the third example of Casella and Strawderman where the unrestricted Bayes rule is minimax, so the Hodges-Lehman’s restriction on point-wise risk is unbinding and the restricted Bayes rule coincides with the unrestricted, but the Huber procedure is inadmissible.
@@
proposed by Efron and Morris. Figure 2 contrasts the Huber and Mallows forms of the Hodges-Lehmann restricted Bayes rule with the unrestricted Bayes rule. We stick to linear shinkage for values of $x$ in the middle but refrain from shrinking at the two tails. Again, the Mallows rule smooths the Huber rule in the center and oscillates around the Huber rule in the tails. Remarkably, Efron and Morris (1971) in their Appendix Figure B’ already illustrates this behavior of the restricted Mallows rule. Extension of this example to James-Stein forms for $G_{0}$ is straightforward although computation of the associated risk is not as simple as shown by Efron and Morris (1972).
@@
$R(1,\delta^{B})=\mathbb{E}_{X\sim\mathcal{N}(1,1)}[\tanh(X)-\theta)^{2}]\approx 0.45<1,$
@@
$\delta^{H}(x)=\begin{cases}x+k&x<-t\\ \delta^{B}(x)&|x|<t\\ x-k&x>t\end{cases}$
with $k=-(\log f_{0})^{\prime}$ and $t$ solves $\int_{-t}^{t}f_{0}(x)dx+\frac{2f_{0}(t)}{K}=(1-\epsilon)^{-1}$, has strictly greater risk for all $\theta\in\Theta=[-1,1]$.
To see this, let $f_{0}(x)=\frac{1}{2}\varphi(x-1)+\frac{1}{2}\varphi(x+1)$, $g^{H}(x)=\delta^{H}(x)-x$ and $g^{B}(x)=\delta^{B}(x)-x=(\log f_{0}(x))^{\prime}$. By Stein’s lemma,
$R(\theta,\delta^{H})$ $-R(\theta,\delta^{B})$
@@
$=-2\mathbb{E}_{\theta}[1\{|X|>t\}(\log f_{0}(X))^{\prime\prime}]$
$\quad+\mathbb{E}_{\theta}[1\{|X|>t\}(k^{2}-((\log f_{0}(X))^{\prime})^{2}]$
The first term is positive because $f_{0}(x)$ is log-concave. The second term is also positive because $k=-(\log f_{0})^{\prime}(t)$ and $k^{2}>((\log f_{0}(X))^{\prime})^{2}$ for $|x|>t$, hence we conclude $\delta^{H}$ is inadmissible. Figure 3 illustrates the pointwise risk functions of the Bayes, Mallows and Huber decision rules on [1,1][-1,1]. Since the Bayes rule and its Mallows modification are identical the two are indistinguishable in the figure.
@@
When the two points of support of $G_{0}$ are more widely separated, for example, $G_{0}=\frac{1}{2}\delta_{2}+\frac{1}{2}\delta_{-2}$, and the support of $\theta$ is the whole real line, the marginal $f_{0}(x)=\int\varphi(x-\theta)dG_{0}(\theta)$ is bimodal, the Huber least favaroable density for $X$ is described in the following proposition.
@@
with $\mathcal{F}_{\epsilon}=\{F:F=(1-\epsilon)F_{0}+\epsilon W\}$ where $F_{0}(x)=\frac{1}{2}\Phi(x+2)+\frac{1}{2}\Phi(x-2)$. The least favorable distribution has its density of the form:
@@ 3.1. Restricted Empirical Bayes rules
In the following theorem, we establish that the excess Bayes risk, the difference between an oracle Mallows rule $\delta^{M}$ with known $G_{0}$ and an empirical Mallows rule $\hat{\delta}^{M}$ with estimated (NPMLE) $G_{0}$ vanishes asymptotically. The proof makes use of machinery from variational analysis and the important feature that the score function of the oracle and EB Mallows least favorable density is uniformly bounded for the oracle and EB Mallows rule. The proof appears in Appendix C.
@@ 4.1. Some Simulation Experience
Appendix A Computation of the Mallows least favorable distribution
As noted by Donoho and Reeves (2013) the Mallows (1978) problem is convex and therefore is easily solved numerically with modern convex optimization software. Our objective is to solve
$\underset{f\in\mathcal{K}}{\min}\int\frac{f^{\prime}(x)^{2}}{f(x)}dx$
with
$\min\sum_{i=1}^{M}w_{i}$
subject to
$u_{i}$ $=f_{i+1}-f_{i},$
$v_{i}$ $=\frac{1}{2}(f_{i+1}+f_{i}),$
$u_{i}^{2}$ $\leq 2v_{i}w_{i}$
Provided that the grids are sufficiently finely spaced interior point optimization in Mosek, ApS (2022), is capable of producing very accurate solutions very efficiently.
In Figure 7 we illustrate a Mallows marginal density, $f^{M}$, its corresponding mixing distribution, $P^{M}$, and plot the log mass of the discrete mass points of $P^{M}$ at their respective locations. On first glance, it seems that the mass points are approximately equally spaced and have mass that declines almost exponentially. However, on closer examination the spacing of the mass points in the right tail are estimated to be: $\{1.96,1.80,1.70,1.61,1.52,1.39,1.29,1.37,1.55,1.70,1.91\}$, which seems sufficiently non-uniform to call the uniform spacing conjecture into question. Donoho et al. (2013) suggest an alternative computational strategy for the Mallows problem using a parametric model that assumes equal spacing of the mass points. The grid for evaluation of $f$ is equally spaced from -30 to 30 with 500 points of evaluation. The grid for evaluation of $P^{M}$ is also equally spaced from -20 to 20 with 4003 points of evaluation. For purposes of illustration the mass at $\theta=0$ is taken to be 0.2.
Figure 8. Mass points of the Mallows contamination distribution $H^{*}$ with $G_{0}=0.5\delta_{-2}+0.5\delta_{2}$. The solid black curve depicts the functional derivative of the objective function of the constrained Mallows’s problem. The vertical green lines depict the location of the mass points of the solution, while their length represents the amount of mass assigned to each. In accordance with the Tukey “hanging rootogram” principle these lengths are rescaled as the square root of the respective masses. The dashed horizontal line represents the bound on the pointwise risk imposed by the Hodges-Lehmann constraint in this case approximately 1.67 induced by the choice of $\epsilon=0.2$.
In the previous example we have taken the initial prior, $G_{0}$ to be point mass one at zero, but there is no obstacle to starting from any other initial prior. To provide some additional intuition about the nature of the Mallows solution for general $G_{0}$, we illustrate in Figure 8 a plot of the functional derivative of the objective evaluated at the solution,
$G^{*}(\theta)=(1-\epsilon)G_{0}(\theta)+\epsilon H^{*}(\theta)$
where $G_{0}(\theta)$ is taken to be an equally weighted mixture of two point masses at -2 and 2. The tangencies in this plot with the horizontal line $R(\delta^{*},\theta)=1+t$ coincide with the location of the mass points of the solution of $H^{*}$ indicated in the plot by the vertical green lines. The bound, $1+t$, is the pointwise risk bound chosen to constrain the modified prior, $G^{*}$ as determined by the choice of $\epsilon$. Since the initial prior places all its mass on the two points $\{-2,2\}$ the Hodges-Lehmann modification hedges this bet by placing a considerable mass at zero and exponentially declining mass at a few points below -2 and above +2. This figure is strongly reminiscent of a similar figure in Lindsay (1995) illustrating the location of mass points of the NPMLE.
Appendix B
Proof. [Proposition 3.1] For all $F\in\mathcal{F}_{\epsilon}$, we can represent its density as $f(x)=(1-\epsilon)f_{0}(x)+\epsilon h(x)$. When $f_{0}$ is symmetric around 0, which is our case here, then the optimal $f^{*}$ will necessarily have $h$ symmetric and hence $f^{*}$ itself symmetric around 0. Let $u(x)=\sqrt{f(x)}$, then $I(F)=4\int(u(x)^{\prime})^{2}dx$ with the constraint $u^{2}(x)\geq(1-\epsilon)f_{0}(x)$ and $\int u^{2}(x)dx=1$. The Lagrangean is then
@@ Appendix A Computation of the Mallows least favorable distribution
with $M:=\{\varphi*P:P\in\mathcal{P}(\mathbb{R})\}$. Let $\hat{f}^{M}_{n}$ be the minimizer of the empirical Bayes version of the Mallows problem
@@
where $\hat{G}_{n}$ denotes the NPMLE of $G$. Denote $f_{0}=\varphi*G_{0}$ and $\hat{f}_{n}=\varphi*\hat{G}_{n}$. The Oracle robust decision rule is $\delta^{M}(x)=x+(\log f^{M}(x))^{\prime}$ and its empirical Bayes counter-part is $\hat{\delta}^{M}(x)=x+(\log\hat{f}_{n}^{M}(x))^{\prime}$. Denote the respective score functions by $g^{M}=(\log f^{M})^{\prime}$ and $\hat{g}^{M}_{n}=(\log\hat{f}_{n}^{M})^{\prime}$. We also know $f^{M}=\varphi*P^{M}$ and $\hat{f}^{M}_{n}=\varphi*P_{n}^{M}$ such that $P_{n}^{M}\in\{G:G=(1-\epsilon)\hat{G}_{n}+\epsilon H\}$ and $P^{M}\in\{G:G=(1-\epsilon)G_{0}+\epsilon H\}$.
@@
Hellinger convergence is established in Jiang and Zhang (2009) under the assumptions they make on $G_{0}$. The strong convergence in $L^{2}$ is a consequence of Lemma C.6. The assumption $I(f_{0})<\infty$ satisfies automatically since $f_{0}\in M$. It provides a bound for $I(f^{M})$ since $f_{0}$ is in the feasible set of Mallows problem, hence $I(f^{M})\leq I(f_{0})<\infty$.
@@
We will show under Assumption C.1, as $n\to\infty$,
@@
Theorem C.1 shows that the first term converges to 0 as $n\to\infty$. The second term
@@
with $C_{1}$ a bounded constant. The last displayed result holds due to Lemma C.9 and the fact that $\|g^{B}\|_{L^{2}(f_{0})}=I(f_{0})<\infty$.
Theorem C.1.
Under Assumption C.1, we have as $n\to\infty$,
@@
    We first establish that $F_{n}$ $\Gamma$-converge to $F$ in $\mathcal{T}$. (Lemma C.3) and $F_{n}$ is equi-coercive (Lemma C.2)
@@
    Then by Lemma C.1, the minimizer of $F$ is unique, we can invoke Corollary 7.24 in Dal Maso (1993) and obtain $\hat{f}^{M}_{n}\rightharpoonup f^{M}$ weakly in $L^{2}$ and $I(\hat{f}^{M}_{n})\to I(f^{M})$.
@@
    Using the fact that $\hat{f}^{M}\in M$ and $f^{M}\in M$, we further show $\hat{P}_{n}^{M}\Rightarrow P^{M}$ and $H(\hat{f}^{M}_{n},f^{M})\to 0$ as $n\to\infty$. (Theorem C.2)
@@
Using the uniform bound (proven in Lemma C.9), we get
@@
Lemma C.3 ($\Gamma$-convergence in weak $L^{2}$).
@@
If $\liminf_{n\to\infty}F_{n}(u_{n})=+\infty$ there is nothing to prove. So we assume $\liminf_{n\to\infty}F_{n}(u_{n})<+\infty$. Set $\ell:=\liminf_{n\to\infty}F_{n}(u_{n})$. Choose indices $n_{k}\uparrow\infty$ such that $F_{n_{k}}(u_{n_{k}})\to\ell$ and $\sup_{k}F_{n_{k}}(u_{n_{k}})<\infty$. Since $u_{n_{k}}\rightharpoonup u$ and $u\in\mathcal{T}$, Lemma C.4 implies the measures $\mu_{k}(dx):=u_{n_{k}}(x)\,dx$ are tight.
For all large $k$, $F_{n_{k}}(u_{n_{k}})<\infty$ implies $u_{n_{k}}\in\mathcal{M}\cap K_{n_{k}}$. Hence there exist $P_{k}\in\mathcal{P}(\mathbb{R})$ such that $u_{n_{k}}=\varphi*P_{k}$ and $P_{k}\geq(1-\epsilon)\hat{G}_{n_{k}}$. By Lemma C.5 and tightness of $\{\mu_{k}\}$, the family $\{P_{k}\}$ is tight. By Prokhorov’s theorem, there exist $k_{j}\uparrow\infty$ and $P\in\mathcal{P}(\mathbb{R})$ such that $P_{k_{j}}\Rightarrow P$. Along the same sub-subsequence, since $\hat{G}_{n}\Rightarrow G_{0}$, $\hat{G}_{n_{k_{j}}}\Rightarrow G_{0}$.
@@
$\int\psi\,u_{n_{k_{j}}}\to\int\psi\,u\qquad\text{and}\qquad\int\psi\,u_{n_{k_{j}}}=\int(\psi*\varphi)(\theta)\,dP_{k_{j}}(\theta)\to\int(\psi*\varphi)(\theta)\,dP(\theta)=\int\psi\,(\varphi*P).$
@@
Finally, boundedness of $F_{n_{k_{j}}}(u_{n_{k_{j}}})$ implies $\sup_{j}I(u_{n_{k_{j}}})<\infty$. By weakly lower semi-continuity of $I(\cdot)$ (Lemma C.7)
@@
which implies $u_{n}\in M\cap K_{n}\subset\mathcal{T}$. Since $\hat{f}_{n}\to f_{0}$ strongly in $L^{2}$ (Lemma C.6), hence $\hat{f}_{n}\rightharpoonup f_{0}$ in $L^{2}$, which implies $u_{n}\rightharpoonup u$ in $L^{2}$.
@@
We have $\hat{G}_{n}\Rightarrow G_{0}$, which implies $\hat{f}_{n}(x)\to f_{0}(x)$ pointwise and $\hat{f}_{n}^{\prime}(x)\to f_{0}^{\prime}(x)$ pointwise, hence $\Psi(u_{n}(x))\to\Psi(u(x))$ pointwise. Moreover apply Lemma C.8, we have
@@
We also have $I(v)<\infty$ since $v\in M$. Apply Pratt’s Lemma (C.10) we obtain $\int\Psi(u_{n})\to\int\Psi(u)$, e.g. $I(u_{n})\to I(u)$.
Theorem C.2.

### 25 Mar 2026 v2

# Robustifying Empirical Bayes

@@ Abstract.
Version: . The authors wish to express their appreciation to Pat Kline, Toru Kitagawa and Peter Bickel for comments on a previous draft. Details on all the computations for the figures and tables are available from https://rkoenker.github.io/www/roger/research/ebayes/ebayes.html.
@@ 1. Introduction
has been proposed as a plug-in estimator for $G$, Jiang and Zhang (2009), Koenker and Mizera (2014) and Soloff et al. (2024). This $G$-modeling strategy – in the terminology of Efron (2019) – performs well in simulations, e.g. Koenker and Mizera (2014), Gu and Koenker (2016), Koenker and Gu (2026), relative to alternatives that attempt to estimate $f_{G}$ directly or that make a priori assumptions about the form of $G$. However, it is obviously subject to the criticism that the Gaussian assumption on the likelihood is quite strong, and priors are never terribly convincing.
@@ 3. Restricted Bayes Solutions
This is the problem posed by Mallows (1978) motivated by robustness considerations for time-series problems with additive outliers. Mallows conjectured that the least favorable $G$ would be discrete, supported on the integers with mass declining exponentially. Bickel (1983) reports a modified conjecture of Donoho that relaxes the spacing of the Mallows mass points, but is otherwise similar. Neither conjecture seems to be strictly correct, but numerical computations confirm the nearly exponential decay of the mass. Bickel and Collins (1983) provide a detailed discussion of the discrete nature of the Mallows solutions based on the analyticity of the objective function. See also Johnstone (1994).
3.1. Computing Mallows’s Least Favorable Distribution
As noted by Bickel and Collins (1983) and Marazzi (1985) the Mallows (1978) problem is convex. Marazzi (1985) suggests a gridding strategy that imposes an exponentially declining mass condition. This produces a remarkably accurate solution for an initial Gaussian prior employing generic optimization software. Using modern convex optimization software, we show that the problem can be efficiently solved numerically for any prior distribution. Our implementation embodied in the function HodgesLehmann in the our REBayes package for the R language employs the Mosek ApS (2022) optimizer and provides a general interface for computing either the Huber or Mallows solutions for an arbitrary initial prior, $G_{0}$.
We now describe our procedure for the simplest (Dirac) initial prior, $G_{0}=\delta_{0}$. Our objective is to solve
$\underset{f\in\mathcal{K}}{\min}\int\frac{f^{\prime}(x)^{2}}{f(x)}dx$
with
$\min\sum_{i=1}^{M}w_{i}$
subject to
$u_{i}$ $=f_{i+1}-f_{i},$
$v_{i}$ $=\frac{1}{2}(f_{i+1}+f_{i}),$
$u_{i}^{2}$ $\leq 2v_{i}w_{i}$
Provided that the grids are sufficiently finely spaced interior point optimization in Mosek is capable of producing very accurate solutions very efficiently.
In Figure 1 we illustrate a Mallows marginal density, $f^{M}$, its corresponding mixing distribution, $P^{M}$, and plot the log mass of the discrete mass points of $P^{M}$ at their respective locations. At first glance, it seems that the mass points are approximately equally spaced and have mass that declines exponentially. However, on closer examination the spacing of the mass points in the right tail are estimated to be: $\{1.96,1.80,1.70,1.61,1.52,1.39,1.29,1.37,1.55,1.70,1.91\}$, which seems sufficiently non-uniform to call the uniform spacing conjecture into question. Donoho et al. (2013) suggest an alternative computational strategy for the Mallows problem using a parametric model that assumes equal spacing of the mass points. The grid for evaluation of $f$ is equally spaced from -30 to 30 with 500 points of evaluation. The grid for evaluation of $P^{M}$ is also equally spaced from -20 to 20 with 4003 points of evaluation. For purposes of illustration the mass at $\theta=0$ is taken to be 0.2.
Figure 2. Mass points of the Mallows contamination distribution $H^{*}$ with $G_{0}=0.5\delta_{-2}+0.5\delta_{2}$ for $\epsilon=0.2$. The solid black curve depicts the pointwise risk function of the decision rule $\delta^{*}$, constructed using Mallows’s least favorable prior $G^{*}$. The vertical green lines depict the location of the mass points of the solution $H^{*}$, while their length represents the amount of mass assigned to each. In accordance with the Tukey “hanging rootogram” principle these lengths are rescaled as the square root of the respective masses. The dashed horizontal line represents the bound on the pointwise risk imposed by the Hodges-Lehmann constraint in this case approximately 1.67 induced by the choice of $\epsilon=0.2$. The constraint is binding at mass points of $H^{*}$.
In the previous example we have taken the initial prior, $G_{0}$ as Dirac, but there is no obstacle to starting from any other initial prior. To provide some additional intuition about the nature of the Mallows solution for general $G_{0}$, we illustrate in Figure 2 a plot of the pointwise risk function $R(\delta^{*},\theta):=\mathbb{E}_{\theta}[(\delta^{*}(X)-\theta)^{2}]$ of the decision rule $\delta^{*}(\cdot)$, constructed as the posterior mean of $\theta$ using the least favorable Mallows prior
$G^{*}(\theta)=(1-\epsilon)G_{0}(\theta)+\epsilon H^{*}(\theta)$
where $G_{0}(\theta)$ is taken to be an equally weighted mixture of two point masses at -2 and 2. The tangencies in this plot with the horizontal dotted line marking out $\sup_{\theta}R(\delta^{*},\theta)=1+t$ coincide with the location of the mass points of the solution of $H^{*}$ indicated in the plot by the vertical green lines. The bound, $1+t$, is the pointwise risk bound chosen to constrain the Hodges-Lehmann restricted Bayes rule, which can be constructed using Mallows’s least favorable prior $G^{*}$. Since the initial prior $G_{0}$ places all its mass on the two points $\{-2,2\}$ the Mallows modification hedges this bet by placing a considerable mass at zero and exponentially declining mass at a few points below -2 and above +2. This figure is strongly reminiscent of Figure 5.5 of Lindsay (1995) illustrating the location of mass points of the NPMLE.
3.2. Some Examples
@@
The Mallows rule has the obvious advantage that it yields a Bayes decision rule while the corresponding Huber procedure does not. This is particularly evident in the third example of Casella and Strawderman where the unrestricted Bayes rule is minimax, so the Hodges-Lehman’s restriction on point-wise risk is unbinding and the restricted Bayes rule coincides with the unrestricted, but the Huber procedure is inadmissible. On the other hand there is something attractive about the Huber rules that it can be shown that they are necessarily monotone. Donoho and Reeves (2013) propose an alternative based on the Huber (1974) spline that minimizes Fisher information over a Kolmogorov neighborhood of the marginal density of $X$ specified by a finite number of evaluations of its quantile function. They then apply the Tweedie formula with the resulting least favorable density. An implementation of this procedure is also included in the REBayes package with the function HuberSpline, although we do not pursue it further in this paper.
@@
proposed by Efron and Morris. Figure 4 contrasts the Huber and Mallows forms of the Hodges-Lehmann restricted Bayes rule with the unrestricted Bayes rule. We stick to linear shrinkage for values of $x$ in the middle but refrain from shrinking at the two tails. Again, the Mallows rule smooths the Huber rule in the center and oscillates around the Huber rule in the tails. Remarkably, Efron and Morris (1971) in their Appendix Figure B’ already illustrates this behavior of the restricted Mallows rule. See also Figure 1 in Marazzi (1985). Extension of this example to James-Stein forms for $G_{0}$ is straightforward although computation of the associated risk is not as simple as shown by Efron and Morris (1972).
@@
$R(\delta^{B},1)=\mathbb{E}_{X\sim\mathcal{N}(1,1)}[\tanh(X)-\theta)^{2}]\approx 0.45<1,$
@@
$\delta^{H}(x)=\begin{cases}x+k&x<-b\\ \delta^{B}(x)&|x|<b\\ x-k&x>b\end{cases}$
with $k=-(\log f_{0})^{\prime}(b)$ and $b$ solves $\int_{-b}^{b}f_{0}(x)dx+\frac{2f_{0}(b)}{k}=(1-\epsilon)^{-1}$, has strictly greater risk for all $\theta\in\Theta=[-1,1]$.
To see this, let $f_{0}(x)=\frac{1}{2}\varphi(x-1)+\frac{1}{2}\varphi(x+1)$, $g^{H}(x)=\delta^{H}(x)-x$ and $g^{B}(x)=\delta^{B}(x)-x=f_{0}^{\prime}(x)/f_{0}(x)$. By Stein’s lemma,
$R(\delta^{H},\theta)$ $-R(\delta^{B},\theta)$
@@
$=-2\mathbb{E}_{\theta}[1\{|X|>b\}(\log f_{0}(X))^{\prime\prime}]$
$\quad+\mathbb{E}_{\theta}[1\{|X|>b\}(k^{2}-((\log f_{0}(X))^{\prime})^{2}]$
The first term is positive because $f_{0}(x)$ is log-concave. The second term is also positive because $k=-(\log f_{0})^{\prime}(b)$ and $k^{2}>((\log f_{0}(x))^{\prime})^{2}$ for $|x|>b$, hence we conclude $\delta^{H}$ is inadmissible. Figure 5 illustrates the pointwise risk functions of the Bayes, Mallows and Huber decision rules on [1,1][-1,1]. Since the Bayes rule and its Mallows modification are identical the two are indistinguishable in the figure.
@@
When the two points of support of $G_{0}$ are more widely separated, for example, $G_{0}=\frac{1}{2}\delta_{2}+\frac{1}{2}\delta_{-2}$, and the support of $\theta$ is the whole real line, the marginal $f_{0}(x)=\int\varphi(x-\theta)dG_{0}(\theta)$ is bimodal, the Huber least favaroable density for $X$ is described in the following proposition whose proof appears in Appendix A.
@@
with $\mathcal{F}_{\epsilon}=\{F:F=(1-\epsilon)F_{0}+\epsilon W\}$ where $F_{0}(x)=\frac{1}{2}\Phi(x+2)+\frac{1}{2}\Phi(x-2)$ and $W$ is arbitrary distribution of $X$. The least favorable distribution has its density of the form:
@@ 3.3. Restricted Empirical Bayes rules
In the following theorem, we establish that the excess Bayes risk, the difference between an oracle Mallows rule $\delta^{M}$ with known $G_{0}$ and an empirical Mallows rule $\hat{\delta}^{M}$ with estimated (NPMLE) $\hat{G}$ vanishes asymptotically. The proof makes use of machinery from variational analysis and the important feature that the score function of the oracle and EB Mallows least favorable density is uniformly bounded. The proof appears in Appendix B.
@@ 3.4. Some simulation experience
For each $G_{0}$ we can evaluate for any $\epsilon$ the corresponding worst case pointwise risk of each rule. The minimax rule achieves a worst case pointwise risk of one, while the Bayes rule typically has unbounded pointwise risk. For the Mallows rule, this can be evaluated by $\mathbb{E}_{\theta^{*}}[(\delta^{M}(X)-\theta^{*})^{2}]$ where $\theta^{*}$ is any mass point of $H^{*}$ as discussed in Section 3.1. For the Huber rule, we can show that for all the $G_{0}$ we considered, $\sup_{\theta}R(\delta^{H},\theta)=1+k_{\epsilon}^{2}$ with $k_{\epsilon}=\sup_{x}|(\log f_{\epsilon}^{H}(X))^{\prime}|$ in which $f_{\epsilon}^{H}$ is the least favorable Huber density for the given $G_{0}$ and $\epsilon$. Details along with some simulation results appear in Appendix C.
@@ 4.1. Some Simulation Experience
Appendix A Proof of Proposition 3.1
Proof. For all $F\in\mathcal{F}_{\epsilon}$, we can represent its density as $f(x)=(1-\epsilon)f_{0}(x)+\epsilon h(x)$. When $f_{0}$ is symmetric around 0, which is our case here, then the optimal $f^{*}$ will necessarily have $h$ symmetric and hence $f^{*}$ itself symmetric around 0. Let $u(x)=\sqrt{f(x)}$, then $I(F)=4\int(u(x)^{\prime})^{2}dx$ with the constraint $u^{2}(x)\geq(1-\epsilon)f_{0}(x)$ and $\int u^{2}(x)dx=1$. The Lagrangean is then
@@
with $M:=\{\varphi*P:P\in\mathcal{P}(\mathbb{R})\}$ where $\mathcal{P}(\mathbb{R})$ is all probability measure supported on $\mathbb{R}$. Let $\hat{f}^{M}_{n}$ be the minimizer of the empirical Bayes version of the Mallows problem
@@
where $\hat{G}_{n}$ denotes the NPMLE of $G$. Denote $f_{0}=\varphi*G_{0}$ and $\hat{f}_{n}=\varphi*\hat{G}_{n}$. The Oracle robust decision rule is $\delta^{M}(x)=x+(\log f^{M}(x))^{\prime}$ and its empirical Bayes counter-part is $\hat{\delta}^{M}(x)=x+(\log\hat{f}_{n}^{M}(x))^{\prime}$. Denote the respective score functions by $g^{M}=(\log f^{M})^{\prime}$ and $\hat{g}^{M}_{n}=(\log\hat{f}_{n}^{M})^{\prime}$. We also know $f^{M}=\varphi*P^{M}$ and $\hat{f}^{M}_{n}=\varphi*P_{n}^{M}$ such that $P_{n}^{M}\in\{P:P=(1-\epsilon)\hat{G}_{n}+\epsilon H\}$ and $P^{M}\in\{P:P=(1-\epsilon)G_{0}+\epsilon H\}$.
@@
Hellinger convergence is established in Jiang and Zhang (2009) under the assumptions they make on $G_{0}$. The strong convergence in $L^{2}$ is a consequence of Lemma B.6. The assumption $I(f_{0})<\infty$ satisfies automatically since $f_{0}\in M$. It provides a bound for $I(f^{M})$ since $f_{0}$ is in the feasible set of Mallows problem, hence $I(f^{M})\leq I(f_{0})<\infty$.
@@
We will show under Assumption B.1, as $n\to\infty$,
@@
Theorem B.1 shows that the first term converges to 0 as $n\to\infty$. The second term
@@
with $C_{1}$ a bounded constant. The last displayed result holds due to Lemma B.9 and the fact that $\|g^{B}\|_{L^{2}(f_{0})}=I(f_{0})<\infty$.
Theorem B.1.
Under Assumption B.1, we have as $n\to\infty$,
@@
    We first establish that $F_{n}$ $\Gamma$-converge to $F$ in $\mathcal{T}$. (Lemma B.3) and $F_{n}$ is equi-coercive (Lemma B.2)
@@
    Then by Lemma B.1, the minimizer of $F$ is unique, we can invoke Corollary 7.24 in Dal Maso (1993) and obtain $\hat{f}^{M}_{n}\rightharpoonup f^{M}$ weakly in $L^{2}$ and $I(\hat{f}^{M}_{n})\to I(f^{M})$.
@@
    Using the fact that $\hat{f}^{M}\in M$ and $f^{M}\in M$, we further show $\hat{P}_{n}^{M}\Rightarrow P^{M}$ and $H(\hat{f}^{M}_{n},f^{M})\to 0$ as $n\to\infty$. (Theorem B.2)
@@
Using the uniform bound (proven in Lemma B.9), we get
@@
Lemma B.3 ($\Gamma$-convergence in weak $L^{2}$).
@@
If $\liminf_{n\to\infty}F_{n}(u_{n})=+\infty$ there is nothing to prove. So we assume that $\liminf_{n\to\infty}F_{n}(u_{n})<+\infty$. Set $\ell:=\liminf_{n\to\infty}F_{n}(u_{n})$. Choose indices $n_{k}\uparrow\infty$ such that $F_{n_{k}}(u_{n_{k}})\to\ell$ and $\sup_{k}F_{n_{k}}(u_{n_{k}})<\infty$. Since $u_{n_{k}}\rightharpoonup u$ and $u\in\mathcal{T}$, Lemma B.4 implies the measures $\mu_{k}(dx):=u_{n_{k}}(x)\,dx$ are tight.
For all large $k$, $F_{n_{k}}(u_{n_{k}})<\infty$ implies $u_{n_{k}}\in\mathcal{M}\cap K_{n_{k}}$. Hence there exist $P_{k}\in\mathcal{P}(\mathbb{R})$ such that $u_{n_{k}}=\varphi*P_{k}$ and $P_{k}\geq(1-\epsilon)\hat{G}_{n_{k}}$. By Lemma B.5 and tightness of $\{\mu_{k}\}$, the family $\{P_{k}\}$ is tight. By Prokhorov’s theorem, there exist $k_{j}\uparrow\infty$ and $P\in\mathcal{P}(\mathbb{R})$ such that $P_{k_{j}}\Rightarrow P$. Along the same sub-subsequence, since $\hat{G}_{n}\Rightarrow G_{0}$, $\hat{G}_{n_{k_{j}}}\Rightarrow G_{0}$.
@@
$\int\psi\,u_{n_{k_{j}}}\to\int\psi\,u\qquad\text{and}\qquad$
$\int\psi\,u_{n_{k_{j}}}=\int(\psi*\varphi)(\theta)\,dP_{k_{j}}(\theta)\to\int(\psi*\varphi)(\theta)\,dP(\theta)=\int\psi\,(\varphi*P).$
@@
Finally, boundedness of $F_{n_{k_{j}}}(u_{n_{k_{j}}})$ implies $\sup_{j}I(u_{n_{k_{j}}})<\infty$. By weakly lower semi-continuity of $I(\cdot)$ (Lemma B.7)
@@
which implies $u_{n}\in M\cap K_{n}\subset\mathcal{T}$. Since $\hat{f}_{n}\to f_{0}$ strongly in $L^{2}$ (Lemma B.6), hence $\hat{f}_{n}\rightharpoonup f_{0}$ in $L^{2}$, which implies $u_{n}\rightharpoonup u$ in $L^{2}$.
@@
We have $\hat{G}_{n}\Rightarrow G_{0}$, which implies $\hat{f}_{n}(x)\to f_{0}(x)$ pointwise and $\hat{f}_{n}^{\prime}(x)\to f_{0}^{\prime}(x)$ pointwise, hence $\Psi(u_{n}(x))\to\Psi(u(x))$ pointwise. Moreover apply Lemma B.8, we have
@@
We also have $I(v)<\infty$ since $v\in M$. Apply Pratt’s Lemma (B.10) we obtain $\int\Psi(u_{n})\to\int\Psi(u)$, e.g. $I(u_{n})\to I(u)$.
Theorem B.2.